# On sporadic symmetry breaking operators for principal series representations of the de Sitter and Lorentz groups

### v1

@@ 3 Applying the F-method
The key tool we use to prove Theorems 1.3 and 1.4 is the F-method. A very powerful machinery developed by T. Kobayashi that reduces the problem of constructing and classifying differential symmetry breaking operators into the computational problem of solving a certain system of ordinary differential equations. We state the main result we use in our setting in Theorem 3.2. This will allow us to derive Theorems 1.3 and 1.4 from Theorem 3.4.
@@
Now, denote by $\mathfrak{g}$ the complexification of the real Lie algebra $\mathfrak{g}({\mathbb{R}})$. Similar notation will be used for other Lie algebras. Let
[…]
$\psi\circ\sigma_{\lambda}^{2N+1}(\ell)=\tau_{m,\nu}(\ell)\circ\text{{Ad}}_{\#}(\ell)\psi,\quad\forall\enspace\ell\in L^{\prime}=M^{\prime}A,$ (3.2)
where the action $\text{{Ad}}_{\#}$ of $L^{\prime}$ on $\text{{Pol}}(\mathfrak{n}_{+})$ is given by
@@
In addition, we set
@@
$\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{{\mathbb{C}}_{m,\nu}}\right)\psi=0,\quad\forall\enspace C\in\mathfrak{n}_{+}^{\prime}.$ (3.3)
@@
$\widehat{d\pi_{\mu}}:\mathfrak{g}\rightarrow\mathcal{D}(\mathfrak{n}_{+})\otimes\text{{End}}(V^{2N+1}_{\lambda}),$
where $\mathcal{D}(\mathfrak{n}_{+})$ denotes the Weyl algebra on $\mathfrak{n}_{+}$. Thus, for any $C\in\mathfrak{n}_{+}$, $\widehat{d\pi_{\mu}}(C)$ defines a differential operator with coefficients on $\text{{End}}(V^{2N+1}_{\lambda})$. In other words, (3.3) defines a differential equation on $\text{{Hom}}_{\mathbb{C}}(V^{2N+1}_{\lambda},{\mathbb{C}}_{m,\nu})\otimes\text{{Pol}}(\mathfrak{n}_{+})$. For details about the definition of $\widehat{d\pi_{\mu}}$, see, for instance [25, Sec. 3] or [30, Sec. 2].
Since $\mathfrak{n}_{+}\simeq{\mathbb{C}}^{3}$ is abelian, we have (see [25, Sec. 4.3])
@@
[…]
By Theorem 3.2 above, in order to determine the space (1.2), it suffices to give a description of
@@
To do so, by (3.2) and (3.3) one may proceed with the following two steps:
  • Step 1. Determine the generators of $\text{{Hom}}_{L^{\prime}}\left(V_{\lambda}^{2N+1},{\mathbb{C}}_{m,\nu}\otimes\text{{Pol}}(\mathfrak{n}_{+})\right)$.
  • Step 2. Solve $\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{{\mathbb{C}}_{m,\nu}}\right)\psi=0,\quad\forall\enspace C\in\mathfrak{n}_{+}^{\prime}$.
@@ 3.1 Description of $\text{{Sol}}(\mathfrak{n}_{+};\sigma_{\lambda}^{2N+1},\tau_{m,\nu})$ for $m>N$
In this section we recall some results proved in [36] that help us to tackle the two steps of the F-method in our setting. In particular, the first step can be achieved by using Proposition 3.3, and the second step can be reduced to solving a system of ODEs by using Theorem 3.5. The complete description of (3.5) is given in Theorem 3.4. We focus mainly on the case $m>N$, since the case $m<-N$ can be deduced from the former one (see Proposition 5.1).
@@
Given $k\in{\mathbb{N}}$, denote by $\mathcal{H}^{k}({\mathbb{C}}^{2})$ the space of harmonic polynomials of degree $k$ on two variables. In other words,
@@
Now, for $|m|\geq N$ and $k\in K_{N,m}:=\{|m-\ell|,|m+\ell|:\ell=0,1,\ldots,N\}$, we define a homomorphism $h_{k}^{\pm}:V^{2N+1}\rightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})$ as follows:
$\begin{array}[]{l l}h_{m+\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m+\ell}({\mathbb{C}}^{2}),&\enspace u_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m+\ell},\\[7.0pt] h_{m-\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m-\ell}({\mathbb{C}}^{2}),&\enspace u_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m-\ell},\\[7.0pt] h_{-m-\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m-\ell}({\mathbb{C}}^{2}),&\enspace u_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m-\ell},\\[7.0pt] h_{-m+\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m+\ell}({\mathbb{C}}^{2}),&\enspace u_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m+\ell}.\end{array}$ (3.7)
Note that the sign index $\pm$ in $h_{k}^{\pm}$ is chosen to coincide with that of m±plus-or-$m\pm\ell$, being $+$ if $m\geq N$ and -- if $m\leq-N$.
@@
We observe that $T_{b}g$ is a homogeneous polynomial on three variables of degree $b$ if $b\in{\mathbb{N}}$ and $g\in\text{{Pol}}_{b}[t]_{\text{{{even}}}}$, where
@@
$\begin{gathered}\text{{Hom}}_{L^{\prime}}\left(V^{2N+1}_{\lambda},{\mathbb{C}}_{m,\nu}\otimes\text{{Pol}}^{a}(\mathfrak{n}_{+})\right)=\text{{span}}_{\mathbb{C}}\{\left(T_{a-k}g_{k}\right)h_{k}^{\text{{sgn}}(m-N)}:g_{k}\in\text{{Pol}}_{a-k}[t]_{\text{{{even}}}},k\in K_{N,m}\},\end{gathered}$
@@
The following three conditions on the quadruple $(\lambda,\nu,N,m)\in{\mathbb{C}}^{2}\times{\mathbb{N}}\times{\mathbb{N}}_{>N}$ are equivalent.
@@
where $a:=\nu-\lambda$ and the polynomials $g_{k}\in\text{{Pol}}_{a-k}[t]_{\text{even}}\enspace(k=m-N,\dots,m+N)$ are given as follows:
@@
3.2 Proof of Theorems 1.3 and 1.4 for $m>N$
We shall discuss the proof of Theorem 3.4 in Section 3.3. By using Theorem 3.4 we first give a proof of Theorems 1.3 and 1.4 for $m>N$. The case $m<-N$ is considered in Section 5.
@@
[…]
@@
$(\text{{Symb}}\otimes\text{{id}})^{-1}\circ\sum_{k=m-N}^{m+N}\left(T_{a-k}g_{k}\right)(\zeta)(\zeta_{1}+i\zeta_{2})^{k}u_{N-m+k}^{\vee},$ (3.14)
@@
×$\enspace\times\left(I_{a-k+2(1-N-\nu+r)}\widetilde{C}_{a-k+2(1-N-\nu+r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3}),$
@@
×$\enspace\times\left(I_{a+k-2(m+N-r)}\widetilde{C}_{a+k-2(m+N-r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3}).$
@@
×$\qquad\times\left(I_{a-k+2(1-N-\nu+r)}\widetilde{C}_{a-k+2(1-N-\nu+r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3})\Big{]},$ $(k=m-N,\ldots,m-1),$
@@
×$\qquad\times\left(I_{a+k-2(m+N-r)}\widetilde{C}_{a+k-2(m+N-r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3})\Big{]},$ $(k=m,\ldots,m+N).$
@@
~$\begin{gathered}\widetilde{{\mathbb{C}}}_{\lambda,\nu}:=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-\Delta_{{\mathbb{R}}^{2}},\frac{\partial}{\partial x_{3}}\right).\end{gathered}$
A change in the indexes given by $d=N-m+k$ gives the following expression of $\mathbb{D}_{\lambda,\nu}^{N,m}$:
@@
$\frac{\partial}{\partial x_{1}}+i\frac{\partial}{\partial x_{2}}=2\frac{\partial}{\partial\overline{z}},\quad\Delta_{{\mathbb{R}}^{2}}=4\frac{\partial^{2}}{\partial z\partial\overline{z}}.$

### v2

@@ 3 Applying the F-method
The key tool we use to prove Theorems 1.3 and 1.5 is the F-method. A very powerful machinery developed by T. Kobayashi that reduces the problem of constructing and classifying differential symmetry breaking operators into the computational problem of solving a certain system of partial differential equations. We state the main result we use in our setting in Theorem 3.2. This will allow us to derive Theorems 1.3 and 1.5 from Theorem 3.4.
@@
Now, denote by $\mathfrak{g}$ the complexification of the real Lie algebra $\mathfrak{g}({\mathbb{R}})$. Similar notation will be used for other Lie algebras. Put $L^{\prime}=M^{\prime}A$ and consider
$\text{{Hom}}_{L^{\prime}}\left(V_{\lambda}^{2N+1},{\mathbb{C}}_{m,\nu}\otimes\text{{Pol}}(\mathfrak{n}_{+})\right),$
where the $L^{\prime}$-action on $\text{{Pol}}(\mathfrak{n}_{+})$ is given by
@@
Now, for $\psi\in\left(\text{{Pol}}(\mathfrak{n}_{+})\otimes(V_{\lambda}^{2N+1})^{\vee}\right)\otimes{\mathbb{C}}_{m,\nu}\simeq\text{{Hom}}_{\mathbb{C}}\left(V_{\lambda}^{2N+1},{\mathbb{C}}_{m,\nu}\otimes\text{{Pol}}(\mathfrak{n}_{+})\right)$, we consider the system
$\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{{\mathbb{C}}_{m,\nu}}\right)\psi=0,\quad\forall\kern 4.62497ptC\in\mathfrak{n}_{+}^{\prime},$ (3.2)
and set
@@
$\widehat{d\pi_{\mu}}:\mathfrak{g}\rightarrow\mathcal{D}(\mathfrak{n}_{+})\otimes\text{{End}}((V^{2N+1}_{\lambda})^{\vee}),$
where $\mathcal{D}(\mathfrak{n}_{+})$ denotes the Weyl algebra on $\mathfrak{n}_{+}$ and $\mu=(\sigma_{\lambda}^{2N+1})^{\vee}\otimes{\mathbb{C}}_{2\rho}$. Thus, for any $C\in\mathfrak{n}_{+}$, $\widehat{d\pi_{\mu}}(C)$ defines a differential operator with coefficients on $\text{{End}}((V^{2N+1}_{\lambda})^{\vee})$. In other words, (3.2) defines a differential equation on $\text{{Pol}}(\mathfrak{n}_{+})\otimes\text{{Hom}}_{\mathbb{C}}(V^{2N+1}_{\lambda},{\mathbb{C}}_{m,\nu})$. For details about the precise definition of $\widehat{d\pi_{\mu}}$, see, for instance [28, Sec. 3] or [33, Sec. 2].
Since $\mathfrak{n}_{+}\simeq{\mathbb{C}}^{3}$ is abelian, we have (see [28, Sec. 4.3])
@@
$\text{{Diff}}_{G^{\prime}}\left(C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1}),C^{\infty}(S^{2},\mathcal{L}_{m,\nu})\right)$$\text{{Sol}}(\mathfrak{n}_{+};\sigma_{\lambda}^{2N+1},\tau_{m,\nu}).$$\scriptstyle{\displaystyle{{}_{\text{{Symb}}\otimes\text{{id}}}}}\scriptstyle{\thicksim}$ (3.3)
By Theorem 3.2 above, in order to determine the space (1.2), it suffices to describe
@@
To do so, by (3.2) one may proceed with the following two steps:
  • Step 1. Determine the generators $\psi$ of $\text{{Hom}}_{L^{\prime}}\left(V_{\lambda}^{2N+1},{\mathbb{C}}_{m,\nu}\otimes\text{{Pol}}(\mathfrak{n}_{+})\right)$.
  • Step 2. Solve $\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{{\mathbb{C}}_{m,\nu}}\right)\psi=0,\quad\forall\kern 4.62497ptC\in\mathfrak{n}_{+}^{\prime}$.
@@ 3.1 Description of $\text{{Sol}}(\mathfrak{n}_{+};\sigma_{\lambda}^{2N+1},\tau_{m,\nu})$ for $m>N$
In this section we recall some results proved in [39] that help us to tackle the two steps of the F-method in our setting. In particular, the first step can be achieved by using Proposition 3.3, and the second step can be reduced to solving a system of ODEs by using Theorem 3.6. The complete description of (3.4) is given in Theorem 3.4. We focus mainly on the case $m>N$, since the case m<Nm<-N can be deduced from the former one (see Proposition 5.1).
@@
Given $k\in{\mathbb{N}}$, denote by $\mathcal{H}^{k}({\mathbb{C}}^{2})$ the space of harmonic homogeneous polynomials of degree $k$ on two variables. In other words,
@@
We recall that this space is given as follows (cf. [17, Thm. 3.1]):
$\mathcal{H}^{k}({\mathbb{C}}^{2})=\begin{cases}{\mathbb{C}},&\text{ if }k=0,\\ {\mathbb{C}}(\zeta_{1}+i\zeta_{2})^{k}\oplus{\mathbb{C}}(\zeta_{1}-i\zeta_{2})^{k},&\text{ if }k>0.\end{cases}$
@@
Now, for $|m|\geq N$ and $k\in K_{N,m}:=\{|m-\ell|,|m+\ell|:\ell=0,1,\ldots,N\}$, we define a linear $SO(2)$-intertwining homomorphism $h_{k}^{\pm}:V^{2N+1}\rightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})$ as follows:
$\begin{array}[]{l l}h_{m+\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m+\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m+\ell},\\[7.0pt] h_{m-\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m-\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m-\ell},\\[7.0pt] h_{-m-\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m-\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m-\ell},\\[7.0pt] h_{-m+\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m+\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m+\ell}.\end{array}$ (3.6)
Note that the sign index $\pm$ in $h_{k}^{\pm}$ is chosen to coincide with that of $m\pm\ell$, being ++ if $m\geq N$ and - if $m\leq-N$.
@@
We observe that $T_{b}g$ is a homogeneous polynomial of degree $b$ on three variables if $b\in{\mathbb{N}}$ and $g\in\text{{Pol}}_{b}[t]_{\text{{{even}}}}$, where
@@
$\begin{gathered}\text{{Hom}}_{L^{\prime}}\left(V^{2N+1}_{\lambda},{\mathbb{C}}_{m,\nu}\otimes\text{{Pol}}^{a}(\mathfrak{n}_{+})\right)=\text{{span}}_{\mathbb{C}}\{\left(T_{a-k}g_{k}\right)h_{k}^{\text{{sgn}}(m)}:g_{k}\in\text{{Pol}}_{a-k}[t]_{\text{{{even}}}},k\in K_{N,m}\},\end{gathered}$
@@
Let $\lambda,\nu\in{\mathbb{C}}$, $N\in{\mathbb{N}}$ and $m\in{\mathbb{Z}}$ such that $m>N$. Then, the following three conditions on the quadruple $(\lambda,\nu,N,m)$ are equivalent.
@@
where $a:=\nu-\lambda$ and the polynomials $g_{k}\in\text{{Pol}}_{a-k}[t]_{\text{even}}\kern 4.62497pt(k=m-N,\dots,m+N)$ are given as follows:
@@
3.2 Proof of Theorems 1.3 and 1.5 for $m>N$
We shall discuss the proof of Theorem 3.4 in Section 3.4. By using Theorem 3.4 we first give a proof of Theorems 1.3 and 1.5 for $m>N$. The case m<Nm<-N is considered in Section 5.
@@
$\text{{Diff}}_{G^{\prime}}\left(C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1}),C^{\infty}(S^{2},\mathcal{L}_{m,\nu})\right)$$\text{{Sol}}(\mathfrak{n}_{+};\sigma_{\lambda}^{2N+1},\tau_{m,\nu}).$$\scriptstyle{\displaystyle{{}_{\text{{Symb}}\otimes\text{{id}}}}}\scriptstyle{\thicksim}$
@@
$(\text{{Symb}}\otimes\text{{id}})^{-1}\left(\sum_{k=m-N}^{m+N}\left(T_{a-k}g_{k}\right)(\zeta)(\zeta_{1}+i\zeta_{2})^{k}u_{N-m+k}^{\vee}\right),$ (3.13)
@@
$\kern 4.62497pt\times\left(I_{a-k+2(1-N-\nu+r)}\widetilde{C}_{a-k+2(1-N-\nu+r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3}),$
@@
$\kern 4.62497pt\times\left(I_{a+k-2(m+N-r)}\widetilde{C}_{a+k-2(m+N-r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3}).$
@@
$\hskip 18.49988pt\times\left(I_{a-k+2(1-N-\nu+r)}\widetilde{C}_{a-k+2(1-N-\nu+r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3})\Big],$ $(k=m-N,\ldots,m-1),$
@@
$\hskip 18.49988pt\times\left(I_{a+k-2(m+N-r)}\widetilde{C}_{a+k-2(m+N-r)}^{\lambda+N-1-r}\right)(-\zeta_{1}^{2}-\zeta_{2}^{2},\zeta_{3})\Big],$ $(k=m,\ldots,m+N).$
@@
$\begin{gathered}\widetilde{{\mathbb{C}}}_{\lambda,\nu}:=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-\Delta_{{\mathbb{R}}^{2}},\frac{\partial}{\partial x_{3}}\right).\end{gathered}$
A change in the indices given by d=Nm+kd=N-m+k gives the following expression of $\mathbb{D}_{\lambda,\nu}^{N,m}$:
@@
$\frac{\partial}{\partial x_{1}}+i\frac{\partial}{\partial x_{2}}=2\frac{\partial}{\partial\overline{z}},\qquad\Delta_{{\mathbb{R}}^{2}}=4\frac{\partial^{2}}{\partial z\partial\overline{z}}.$
Remark 3.5.
The results obtained in [38] for the case $|m|>N=1$ coincide with the ones obtained in this paper. In fact, the condition on the parameters $(\lambda,\nu)$ in [38, Thm. 1.2] clearly coincides with that given in Theorem 1.3. As for the generator of (1.2), in [38] it is given as follows for $\lambda\in{\mathbb{Z}}_{\leq 1-|m|}$ and $\nu=0,1,2$.
[…]
$2^{2\nu}(-1)^{\nu+1}\frac{\Gamma\left(\lambda+\left[\frac{-\lambda-\nu-m+2}{2}\right]\right)}{\Gamma(\lambda+\left[\frac{\nu-\lambda-m-1}{2}\right])}\widetilde{{\mathbb{C}}}_{\lambda+1,2-\nu-m}\frac{\partial^{2\nu+m-1}}{\partial z^{\nu}\partial\overline{z}^{\nu+m-1}}\otimes u_{0}^{\vee}\\ +\left((m(1-\nu)-\lambda+2)\widetilde{{\mathbb{C}}}_{\lambda,\nu-m}-2\gamma(\lambda-1,\nu-\lambda-m)\widetilde{{\mathbb{C}}}_{\lambda+1,\nu-m}\frac{\partial}{\partial x_{3}}\right)\frac{\partial^{m}}{\partial\overline{z}^{m}}\otimes u_{1}^{\vee}\\ -4\gamma(\lambda-1,\nu-\lambda-m)\widetilde{{\mathbb{C}}}_{\lambda+1,\nu-m}\frac{\partial^{m+1}}{\partial\overline{z}^{m+1}}\otimes u_{2}^{\vee},$ (3.15)
$4\gamma(\lambda-1,\nu-\lambda+m)\widetilde{{\mathbb{C}}}_{\lambda+1,\nu+m}\frac{\partial^{-m+1}}{\partial z^{-m+1}}\otimes u_{0}^{\vee}\\ +\left((m(1-\nu)+\lambda-2)\widetilde{{\mathbb{C}}}_{\lambda,\nu+m}+2\gamma(\lambda-1,\nu-\lambda+m)\widetilde{{\mathbb{C}}}_{\lambda+1,\nu+m}\frac{\partial}{\partial x_{3}}\right)\frac{\partial^{-m}}{\partial z^{-m}}\otimes u_{1}^{\vee}\\ +2^{2\nu}(-1)^{\nu+1}\frac{\Gamma\left(\lambda+\left[\frac{-\lambda-\nu+m+2}{2}\right]\right)}{\Gamma(\lambda+\left[\frac{\nu-\lambda+m-1}{2}\right])}\widetilde{{\mathbb{C}}}_{\lambda+1,2-\nu+m}\frac{\partial^{2\nu-m-1}}{\partial z^{\nu-m-1}\partial\overline{z}^{\nu}}\otimes u_{2}^{\vee}.$ (3.16)
Here, $\gamma(\mu,\ell)$ denotes the constant defined in (8.11).
$\mathbb{D}_{\lambda,\nu}^{m}=2\mathbb{D}_{\lambda,\nu}^{1,m}.$ (3.17)
In fact, by (1.5), $\mathbb{D}_{\lambda,\nu}^{1,m}$ amounts to
On the other hand, from the three-term relation [38, Eq. (9.5)], we have
$=\kern 4.62497pt$ $(m+1)(2-\nu)\widetilde{{\mathbb{C}}}_{\lambda,\nu-m}+8\widetilde{{\mathbb{C}}}_{\lambda+1,\nu-m-1}\frac{\partial^{2}}{\partial z\partial\overline{z}}.$
And by (1.4), the following holds:
$2B(1,0)$ $=4,$
$\frac{1}{2}B(1,1)$ $=\frac{1}{2}(m+1)(2-\nu),$
$B(2,0)$ $=-2\gamma(\lambda-1,\nu-\lambda-m).$
Now, (3.17) clearly follows.
3.3 Proof of Corollary 1.4
Proof of Corollary 1.4.
Suppose that there exist a non-zero differential symmetry breaking operator $\mathbb{D}$ in (1.2). Consider the representations $\mathcal{H}^{N}({\mathbb{C}}^{3})\in\widehat{O(3)}$ and $\mathcal{H}^{|m|}({\mathbb{C}}^{2})\in\widehat{O(2)}$, and following the notation given in [31, Sec. 2.3], consider the principal series representations
$I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)\qquad\text{and}\qquad J(\mathcal{H}^{|m|}({\mathbb{C}}^{2}),\nu)$
of $O(4,1)$ and $O(3,1)$, respectively. Now, since
$\mathcal{H}^{N}({\mathbb{C}}^{3})\big\rvert_{SO(3)}\simeq V^{2N+1},\text{ as }SO(3)\text{-modules},$
$\mathcal{H}^{|m|}({\mathbb{C}}^{2})\big\rvert_{SO(2)}\simeq{\mathbb{C}}_{m}\oplus{\mathbb{C}}_{-m},\text{ as }SO(2)\text{-modules},$
we clearly have
$I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)\big\rvert_{SO_{0}(4,1)}\simeq C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1}),\text{ as }SO_{0}(4,1)\text{-modules},$ (3.18)
$J(\mathcal{H}^{|m|}({\mathbb{C}}^{2}),\nu)\big\rvert_{SO_{0}(3,1)}\simeq C^{\infty}(S^{2},\mathcal{L}_{m,\nu})\oplus C^{\infty}(S^{2},\mathcal{L}_{-m,\nu}),\text{ as }SO_{0}(3,1)\text{-modules}.$ (3.19)
By the irreducibility criterion [31, Thm. 4.15] and by [31, Ex. 2.8], $I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)$ is irreducible as an $O(4,1)$-module if and only if $\lambda\in\{1-N,2+N\}$. On the other hand, by Theorem 1.3 we must have $\lambda\in{\mathbb{Z}}_{\leq 1-|m|}$. In particular, $\lambda\not\in\{1-N,2+N\}$ as $|m|>N$. Therefore, $I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)$ is reducible as an $O(4,1)$-module, and by [31, Lem. 15.3] and (3.18), $C^{\infty}(S^{3},V_{\lambda}^{2N+1})$ is also reducible as an $SO_{0}(4,1)$-module. This proves (1).
In order to prove (2), we follow a similar argument. By [31, Thm. 4.15] and [31, Ex. 2.8], $J(\mathcal{H}^{|m|}({\mathbb{C}}^{2}),\nu)$ is irreducible as an $O(3,1)$-module if and only if $\nu\in\{1-|m|,1+|m|\}\cup\{\nu\in{\mathbb{Z}}:0<|\nu-1|<|m|\}=[1-|m|,1+|m|]\cap{\mathbb{Z}}$. On the other hand, by Theorem 1.3 we must have $\nu\in[1-N,N+1]\cap{\mathbb{Z}}$. However, since $|m|>N$, we have $1-N,N+1]\subset[1-|m|,1+|m|]$. Hence, $J(\mathcal{H}^{|m|}({\mathbb{C}}^{2}),\nu)$ is irreducible as an $O(3,1)$-module, and by [31, Lem. 15.3] and (3.19), $C^{\infty}(S^{2},\mathcal{L}_{\pm m,\nu})$ is also irreducible as an $SO_{0}(3,1)$-module. In particular, $\overline{\text{{im}}(\mathbb{D})}=C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$. ∎